Grade 10Algebra

Let n>1 be an integer and f(x)=xn+5xn-1+3.Show that f(x) is irreducible over Z.

Ankan Sen
8 Years agoGrade 10

1 Answer

shubham
8 Years ago
let a be integer, and a zeo of above expression , we apply contradiction
if
 a=even
an=even=2p
an-1=2q
the expression=xn+5xn-1+3=2p+10q+3=which cannot be equal to 0 as even+even+odd=odd
if 
 a=odd
an=2p+1
an-1=2q+1
expression reduces to
2p+10q+9=which cannot be equal to zero = as e+e+o=o;